# Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface

### v1

@@ Abstract
Having ascertained the existence of rough and crumpled phases in 2D, we now speculate on the global fixed point structure including the strong coupling fixed points, which are inaccessible to perturbative RG; see Fig. 2. The RG flow lines that flow out of the fixed point (0,0) in the unstable region of the phase space should flow to one of these fixed points. To analyze these fixed points, we are first guided by the expectation that there should be a strong coupling fixed point at $g\neq 0,g_{2}=0$, which is the expected 2D KPZ string coupling fixed point and should govern the scaling laws for a 2D KPZ surface (which is rough). This is marked by a red circle (the other red circle at the origin is another fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction). In the present model, this fixed point is expected to be unstable against perturbation by a finite $\gamma$ or $g_{2}$. On one side of the $g$-axis, there should be a crumpled phase, which we expect to be characterized by a stable crumpled phase fixed point, marked by a blue square in Fig. 2. On the other side of the $g$-axis, there should be a rough phase. Considering the MCT prediction of nonuniversal scaling in the rough phase, we speculate a fixed line marked by broken line in Fig. 2, corresponding to the rough phase. On the side of the $g$-axis with a crumpling fixed, we also expect to have a stable rough phase fixed line for higher negative values of $g_{2}$ (or higher negative values of $\gamma$), giving nonuniversal scaling. Finally, we generally expect that for larger noises, i.e., with larger values of $D$, for which both $g$ and $g_{2}$ grow, the surface should be less and less ordered. This line of reasoning suggests that there should be a crumpling fixed point on the $g$-axis beyond the putative 2D KPZ rough phase fixed point, as well as on the $g_{2}$ axis (marked by a green triangle), indicating eventual crumpling of the surface for high enough $g$ or $g_{2}$. These are basically an “Occam”-razor style arguments: Fig. 2 has the simplest topology that naturally reduces to the known RG flow lines for small $g,\,g_{2}$, as shown in Fig. 1(a). At the same time, it gives the putative global flow lines, allowing for a transition to presumed rough and crumpled phases; see Ref. Banerjee et al. (2019); Mukherjee and Basu (2022a, b) for similar Occam’s razor arguments in different contexts.
@@
We now consider the model in higher dimensions: $d>2$. In $d=2+\epsilon>2$, there are additional fixed points of the RG flows, other than (0,0) (stable fixed point) in the $g$-$g_{2}$ plane. In the accessible region i.e. $\tilde{A}(\gamma)>0$, (0,0) is the only stable fixed point and the long-wavelength scaling behavior matches that of the Edwards-Wilkinson (EW) equation, characterized by $z=2,\,\chi=(2-d)/2$ Jana et al. (2024). In the perturbatively inaccessible strong coupling phase, i.e., $\tilde{A}(\gamma)<0$, from an RG analysis Ref. Jana et al. (2024) argued a roughening transition with an associated $\chi=-\epsilon\,\tilde{C}(\gamma)/{\tilde{A}(\gamma)}$; $\tilde{C}(\gamma)\equiv\gamma^{2}/2+5\gamma/8$. For any $\epsilon>0$, at the roughening transition $\chi$ can grow arbitrarily large, if $\tilde{A}(\gamma)$ becomes small enough (remaining negative). As $\chi$ exceeds unity, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface. (We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq. (2) to breakdown.) Setting $\chi=1$, we get a limit on $\gamma$, up to which this roughening transition between a smooth phase and an algebraically rough phase can be observed. Beyond this threshold value of $\gamma$ there is a range of $\gamma$ bounded by another threshold such that in the entire range $\tilde{A}(\gamma)$ is small enough to make $\chi>1$, suggesting that the roughening transition in this range of $\gamma$ is actually between a smooth phase and a crumpled phase (and not a rough phase)! Overall thus, similar to our discussions for 2D, we expect the strong coupling phase to consist of a rough phase with $10<\chi<$ 1 and a crumpled phase with $\chi>1$. Our further theoretical analysis below substantiate this expectation.
@@
We now briefly discuss the MCT predictions for $d>2$. Our RG calculations already predicted a smooth phase and a roughening transition presumably to a strong coupling phase. The latter however cannot be accessed in the RG calculations. Assuming $\chi>0$, we apply MCT to extract the scaling behaviour in the strong coupling phase. Focusing on $d=3$, we get $A=36\gamma^{2}+48\gamma+15$, $B=54\gamma^{2}+78\gamma-2.5$ and $C=16\gamma^{2}+24\gamma$, giving $\chi\equiv\chi(\gamma)$. Thus, both $\chi$ and $z=2-\chi$ are nonuniversal, as they are in 2D. Secondly, by varying $\gamma$, one can make $\chi$ bigger, eventually exceeding unity. As in 2D, $\chi>1$ should imply crumpling of the surface, for which our theory breaks down. We therefore find that the strong coupling phase in 3D is similar to 2D. Similar analysis indicates analogous behavior in the strong coupling phase for any $d>2$.
@@
We have thus studied the strong coupling phases of a generalized KPZ equation, in which local and nonlocal nonlinear effects compete (in the scaling sense). Our results from the RG and bare perturbation theory in 2D suggest that the strong coupling region in the parameter space the strong coupling phase actually is made of two distinct phases, one an algebraically rough phase with $10<\chi<$ 1 corresponding to orientational LRO but positional SRO, and another with $\chi>1$ implying a crumpled phase having both positional and orientational SRO. These are corroborated by our one-loop MCT calculations. While the crumpled phase cannot be systematically explored by our theory, our MCT predicts that the algebraic rough phase is characterized by nonuniversal scaling exponents parametrized by $\gamma$, similar to the nonuniversal exponents in the logarithmically rough phase by the RG calculations. Similar behavior including a crumpled phase and a rough phase with nonuniversal scaling exponents are predicted by MCT in 3D. The results in Ref. Jana et al. (2024) showed that chirality has the effect of suppressing the nonlinear instabilities of the RG flow. We thus expect that chiral effects together with the nonlocality should be able to suppress the crumpling of the membrane, stabilising orientational order. We note that the nonuniversal scaling exponents are a crucial outcome of the lack of vertex renormalization in one-loop MCT. We are unable to speculate whether this result protected by any “hidden” symmetry not known at present, or a fortuitous result, or will not hold at higher order perturbation theory. Thus, it would be interesting to verify our results numerically, e.g., by using pseudo-spectral methods Basu and Frey (2009a); Basu et al. (1998), or by nonperturbative methods Canet et al. (2011a, b, 2010). Nonetheless, we expect our predictions on the existence of a crumpled phase to hold true.
@@
Acknowledgement:- A.B. thanks the SERB, DST (India) for partial financial support through the TARE scheme [file no.: TAR/2021/000170] and Alexander von Humboldt Stiftung, Germany for partial financial support through the Research Group Linkage Programme (2024).
Appendix A Renormalization group calculations
@@
We revisit and reanalyse the renormalization group calculations on the generalized KPZ equation (2) of the main text. We follow Jana et al. (2024). We first give the path integral over $h({\bf r},t)$ and its dynamic conjugate field $\hat{h}({\bf r},t)$ Bausch et al. (1976); Täuber (2014) that is equivalent to and constructed from Eq. (2) of the main text together with the noise variance. The generating functional corresponding to Eq. (2) of the main text is given by Bausch et al. (1976); Täuber (2014)
@@
$S=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)\bigg{\}}.$ (17)
Using the one loop diagrams from the supplemental material of Ref. Jana et al. (2024) we obtain renormalized $D$ ($D^{<}$) and renormalized $\nu$ ($\nu^{<}$):
@@
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, where $S_{d}$ is the surface area of a $d$-dimensional sphere and $\gamma=\frac{\lambda_{1}}{\lambda}$ as defined in the main text. Using above equations and Setting $d=2$ in the above integrals, with $b=e^{\delta l}\approx 1+\delta l$ ($l$ is the RG time) and defining dimensionless coupling constant by $g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d}$ as in the main text, we obtain the RG flow equations for $D$ and $\nu$:
@@
Also defining another dimensionless coupling constant, $g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}$ and using above equations we can write the flow equations of $g$ and $g_{2}$:
@@
$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)].$ (23)
@@
Appendix B MCT Calculations
We start with the general definitions of the correlation $C({{\bf k}},\omega)$ and response functions $G({{\bf k}},\omega)$
@@ Appendix B MCT Calculations
The response and correlation functions are assumed to have the following scaling form
@@
where $z$ is the dynamic exponent and $\chi$ is the roughness exponent. The Dyson’s equation for the self energy in the scaling limit is
@@
Here $\Gamma$ is basically renormalised or effective $\nu$. The correlation function in Lorentzian approximation can be written as
@@
Figure 3(a) shows the one-loop diagrammatic correction to the self energy at ${\cal O}(\lambda)^{2}$ that comes with a symmetry factor of $8$ and contributes
$\Bigl{(}\frac{\lambda}{2}\Bigl{)}^{2}\times\frac{1}{2!}\times 8\int_{{\bf q},\Omega}{\bf q\cdot(k-q)(q\cdot k)}C({\bf q},\Omega)G({\bf k-q},-\Omega)=\lambda^{2}\int_{{\bf q},\Omega}\frac{{\bf(k-q)\cdot q\times(k\cdot q)}\times 2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})(i\Omega+\Gamma({\bf k-q})^{z})}$ (32)
After performing the $\Omega$ integral and making ${\bf q}\rightarrow{\bf(q+\frac{k}{2})}$ above integral becomes,
$I$ $=\frac{\lambda^{2}D}{\Gamma}\int_{\bf q}\Bigl{(}{\bf q+\frac{k}{2}}\Bigl{)}^{-2\chi-d+z}{\bf\biggl{[}\Bigl{(}\frac{k}{2}-q\Bigl{)}\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}\frac{{\bf\biggl{[}k\cdot\Bigl{(}q+\frac{k}{2}\Bigl{)}\biggl{]}}}{{\bf\Bigl{(}q+\frac{k}{2}\Bigl{)}}^{z}+{\bf\Bigl{|}q-\frac{k}{2}\Bigl{|}}^{z}}$
$=\frac{\lambda^{2}D}{2\Gamma}\frac{2\chi+d}{2}\int_{\bf q}\frac{q^{2}({\bf k\cdot q})q^{-2\chi-d}}{q^{z}}\frac{({\bf k\cdot q})}{q^{2}}-\frac{\lambda^{2}D}{2\Gamma}k^{2}\int_{\bf q}\frac{q^{2}}{2q^{z}}q^{-2\chi-d}$ (33)
@@
$\frac{\lambda^{2}D}{2\Gamma}\times\frac{2\chi+d}{2}\times\frac{k_{d}}{d}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
@@
$\frac{\lambda^{2}D}{2\Gamma}\times\frac{k_{d}}{2}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
@@
Figure 3(b) shows the one-loop contribution to the self energy at ${\cal O}(\lambda_{1})^{2}$ with a symmetry factor of $8$ and contributes
$I$ $=\frac{\lambda_{1}^{2}}{2!}\times 8\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}q_{i}(k-q)_{j}\frac{(q-k)_{m}(q-k)_{n}}{(q-k)^{2}}q_{m}k_{n}C({\bf q},\Omega)\times G({\bf k-q},-\Omega)$
$=\frac{4\lambda_{1}^{2}k_{i}k_{j}k_{n}}{k^{2}}\int_{{\bf q},\Omega}q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}\frac{1}{{\bf(q-k)^{2}}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})}\times\frac{1}{i\Omega+\Gamma{\bf(k-q)}^{z}}$ (35)
After performing the $\Omega$ integral, we get
@@
Then shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, above integral becomes
$\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q-\frac{k}{2})_{m}(q-\frac{k}{2})_{n}(q+\frac{k}{2})_{m}}{{\bf(q-\frac{k}{2})}^{2}({\bf|q+\frac{k}{2}|}^{z}+{\bf|q-\frac{k}{2}|}^{z})}{\bf(q+\frac{k}{2})}^{-2\chi-d}.$ (37)
@@
$I_{2}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{j}q_{m}q_{n}q^{-2\chi-d}}{q^{2}\times 2q^{z}}\times\frac{2\chi+d}{2}=\frac{3\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{2\chi+d}{\chi}k_{d}k^{-\chi+2},$ (39)
where we have used the following identity Yakhot and Orzag (1986),
@@
Then, $I_{3}$ is evaluated by using the well-known relation Yakhot and Orzag (1986)
@@
Adding $I_{1}$, $I_{2}$ and $I_{3}$ we get contribution to zero frequency self energy from vertices $\lambda_{1}$-$\lambda_{1}$
@@
Figure 3(c) gives one of the two one-loop corrections to the self energy at ${\cal O}(\lambda_{1}\lambda)$,with a contribution
@@
After performing the $\Omega$ integral and shifting ${\bf q}\rightarrow{\bf q+\frac{K}{2}}$, we get
$\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}\frac{(q+\frac{k}{2})^{-2\chi-d}}{{\bf(q+\frac{k}{2})}^{z}+{\bf|\frac{k}{2}-q|}^{z}}.$ (45)
@@
In evaluation of these integrals we used identity (40) and (42). After some simplifications, the above integral gives
@@
After performing the $\Omega$ integral above integral is,

### v3

@@ Abstract
Having ascertained the existence of rough and crumpled phases in 2D, we now speculate on the global fixed point structure, including the strong coupling fixed points, which are inaccessible to perturbative RG; see Fig. 2. The RG flow lines that flow out of the fixed point (0,0) in the unstable region of the phase space should flow to one of these fixed points. To analyze these fixed points, we are first guided by the expectation that there should be a strong coupling fixed point at $g\neq 0,g_{2}=0$, which is the expected 2D KPZ strong coupling fixed point and should govern the scaling laws for a 2D KPZ surface (which is rough). This is marked by a red circle (the other red circle at the origin is another fixed point that is unstable along the $g$ direction but stable along the $g_{2}$ direction). In the present model, this fixed point is expected to be unstable against perturbation by a finite $\gamma$ or $g_{2}$. On one side of the $g$ axis, there should be a crumpled phase, which we expect to be characterized by a stable crumpled phase fixed point, marked by a blue square in Fig. 2. On both sides of the $g$ axis, there should be a rough phase enclosing the $g$ axis. Considering the MCT prediction of nonuniversal scaling in the rough phase, we speculate a fixed line in each of these regions marked by broken line in Fig. 2, corresponding to the rough phase. On the side of the $g$ axis with a crumpling fixed point, we also expect to have another region with a stable rough phase fixed line for higher negative values of $g_{2}$ (or higher negative values of $\gamma$), giving nonuniversal scaling. Finally, we generally expect that for larger noises, i.e., with larger values of $D$, for which both $g$ and $g_{2}$ grow, the surface should be less and less ordered. This line of reasoning suggests that there should be two roughening fixed points on the $g_{2}$ axis on either side of the $g$ axis (marked by inverted triangles), describing roughening of logarithmically rough crumpling fixed points on the $g$ axis beyond the putative 2D KPZ rough phase fixed point, as well as on the $g_{2}$ axis (marked by green triangles), indicating eventual crumpling of the surface for high enough $g$ or $g_{2}$. These are basically “Occam’s razor” style arguments: Fig. 2 has the simplest topology that naturally reduces to the known RG flow lines for small $g,\,g_{2}$, as shown in Fig. 1(a). At the same time, it gives the putative global flow lines, allowing for a transition to presumed rough and crumpled phases; see Refs. Banerjee et al. (2019); Mukherjee and Basu (2022a, b) for similar Occam’s razor arguments in different contexts.
@@
We now consider the model in higher dimensions $d>2$. In $d=2+\epsilon>2$, there are additional fixed points of the RG flows other than (0,0) (stable fixed point) in the $g$-$g_{2}$ plane. In the accessible region, i.e., $\tilde{A}(\gamma)>0$, (0,0) is the only stable fixed point and the long-wavelength scaling behavior is identical to that of the Edwards-Wilkinson (EW) equation, characterized by $z=2,\,\chi=(2-d)/2$ Jana et al. (2024). In the perturbatively inaccessible strong coupling phase, i.e., $\tilde{A}(\gamma)<0$, from an RG analysis Ref. Jana et al. (2024) argued a roughening transition with an associated $\chi=-\epsilon\,\tilde{C}(\gamma)/{\tilde{A}(\gamma)}$; $\tilde{C}(\gamma)\equiv\gamma^{2}/2+5\gamma/8$. For any $\epsilon>0$, at the roughening transition $\chi$ can grow arbitrarily large, if $\tilde{A}(\gamma)$ becomes small enough (remaining negative). As $\chi$ exceeds unity, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface. [As before, it suggests that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in the RG sense), making our theory given by Eq. (2) to breakdown.] Setting $\chi=1$, we get a limit on $\gamma$, up to which this roughening transition between a smooth phase and an algebraically rough phase can be observed. Beyond this threshold value of $\gamma$, there is a range of $\gamma$ bounded by another threshold such that in the entire range $\tilde{A}(\gamma)$ is small enough to make $\chi>1$, suggesting that the roughening transition in this range of $\gamma$ is actually between a smooth phase and a crumpled phase (and not a rough phase)! Overall thus, similar to our discussions for 2D, we expect the strong coupling phase to consist of a rough phase with $0<\chi<1$ and a crumpled phase with $\chi>1$. See Fig. 3.
Figure 3: Schematic phase diagram together with RG flow lines in the $g$-$g_{2}$ plane for $d>2$. The blue square at the origin (0,0) is a stable fixed point that controls the scaling of the smooth phase. The red circle is an unstable fixed point found in the KPZ equation at $d>2$, which demarcates the strong coupling rough phase and the smooth phase. The solid curve lines delineate the strong coupling phases and the smooth phase. Within the strong coupling phase regions, the small region bounded between the broken lines, drawn schematically, correspond to the crumpled phase, the rest being the usual rough phase. Arrows indicate the RG flow directions.
@@
We now briefly discuss the MCT predictions for $d>2$ for the strong coupling phases in Fig. 3. Our RG calculations already predicted a smooth phase and a roughening transition presumably to a strong coupling phase. The latter, however, cannot be accessed in the RG calculations. Assuming $\chi>0$, we apply MCT to extract the scaling behavior in the strong coupling phase. Focusing on $d=3$, we get $A=36\gamma^{2}+48\gamma+15$, $B=54\gamma^{2}+78\gamma-2.5$, and $C=16\gamma^{2}+24\gamma$, giving $\chi\equiv\chi(\gamma)$. Thus, both $\chi$ and $z=2-\chi$ are nonuniversal, as they are in 2D. Secondly, by varying $\gamma$, one can make $\chi$ bigger, eventually exceeding unity. As in 2D, $\chi>1$ should imply crumpling of the surface, for which our theory breaks down. We therefore find that the strong coupling phase in 3D is similar to 2D. Similar analysis indicates analogous behavior in the strong coupling phase for any $d>2$.
@@
We have thus studied the strong coupling phases of a generalized KPZ equation in which local and nonlocal nonlinear effects compete (in the scaling sense). Our results from the RG and bare perturbation theory in 2D suggest that the strong coupling phase region in the parameter space actually is made of two distinct phases, one an algebraically rough phase, with $0<\chi<1$ corresponding to orientational LRO but positional SRO, and another with $\chi>1$, implying a crumpled phase having both positional and orientational SRO. These are corroborated by our one-loop MCT calculations. The crumpled phase, however, cannot be systematically explored by our theory. Our MCT predicts that the algebraic rough phase is characterized by nonuniversal scaling exponents parametrized by $\gamma$, similar to the nonuniversal exponents in the logarithmically rough phase predicted by the RG calculations. Similar behavior including a crumpled phase and a rough phase with nonuniversal scaling exponents are predicted by MCT in 3D. The results in Ref. Jana et al. (2024) showed that chirality has the effect of suppressing the nonlinear instabilities of the RG flow. We thus expect that chiral effects together with the nonlocality should be able to suppress the crumpling of the membrane, stabilizing orientational order. We note that the nonuniversal scaling exponents found in the MCT calculations are a crucial outcome of the lack of vertex renormalization in one-loop MCT. We are unable to speculate whether this result is protected by any “hidden” symmetry not known at present, or a fortuitous result, or will not hold at higher order perturbation theory. Indeed, if there are unequal infinite renormalizations of $\lambda,\,\lambda_{1}$ at some higher loop order, their ratio $\gamma$ will acquire nontrivial fixed point value(s), instead of being marginal. In this case, a rough phase would be characterized by isolated fixed point(s), which should replace the fixed lines in the Occam’s razor global RG flow diagram (2). Thus, it would be interesting to verify our results numerically, e.g., by using pseudospectral methods Basu and Frey (2009); Basu et al. (1998) or by nonperturbative methods Canet et al. (2010, 2011a, 2011b). However, we expect our predictions on the existence of a crumpled phase to hold true. Our theory should apply to nonequilibrium surfaces with other fast relaxing active degrees of freedom living on it. We look for future experimental endeavors along this direction.
@@
Acknowledgement. A.B. thanks the Alexander von Humboldt Stiftung, Germany for partial financial support through the Research Group Linkage Programme (2024). The authors thank T. Halpin-Healy for various helpful suggestions.
Supplemental Material for “Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface”
I Galilean invariance
We rewrite the generalized KPZ equation Eq. (2) in the main text, which was introduced as a conceptual model incorporating competing local and nonlocal nonlinear interactions,
$\frac{\partial h}{\partial t}$ $=\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}+\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta,$ (16)
where $Q_{ij}({\bf x})\equiv\nabla_{i}\nabla_{j}/\nabla^{2}$ is the longitudinal projection operator and takes the form $k_{i}k_{j}/k^{2}$ in the Fourier space; $\bf k$ is a Fourier wavevector. Under the transformation $x_{i}^{\prime}\rightarrow x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}\rightarrow t$, where $i$ and $j$ can take values 111 or 222 we obtain,
$\frac{\partial}{\partial t^{\prime}}\rightarrow\frac{\partial}{\partial t}+(\lambda+2\lambda_{1})\textbf{c}\cdot\bm{\nabla};\quad\bm{\nabla}^{\prime}\rightarrow\bm{\nabla}.$ (17)
Using these we show below that if the height function $h({\bf x},t)$ transforms as $h^{\prime}({\bf x}^{\prime},t^{\prime})\rightarrow h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, then Eq. (16) remains invariant. Considering RHS of Eq. (16), the first term transforms as
$\nu\nabla^{\prime 2}h^{\prime}\rightarrow\nu\nabla^{2}h.$ (18)
The second term transforms as
$\frac{\lambda}{2}(\bm{\nabla}^{\prime}h^{\prime})^{2}\rightarrow\frac{\lambda}{2}(\bm{\nabla}h)^{2}+\lambda\bm{\nabla}h\cdot\textbf{c}+\frac{\lambda}{2}c^{2}.$ (19)
The third term transforms as
$\lambda_{1}\frac{\nabla^{\prime}_{i}\nabla^{\prime}_{j}}{\nabla^{\prime 2}}(\nabla^{\prime}_{i}h^{\prime}\nabla^{\prime}_{j}h^{\prime})\rightarrow\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)+2\lambda_{1}\textbf{c}\cdot\bm{\nabla}h.$ (20)
Noise term remains invariant. Similarly LHS of Eq. (16) transforms as
$\frac{\partial h^{\prime}}{\partial t^{\prime}}\rightarrow\frac{\partial h}{\partial t}+(\lambda+2\lambda_{1})\textbf{c}\cdot\bm{\nabla}h+(\lambda+2\lambda_{1})c^{2}.$ (21)
From the above equations we see that Eq. (16) is invariant.
II Renormalization group calculations
@@
We revisit and reanalyze the renormalization group calculations on the generalized KPZ equation (2) of the main text. We follow Jana et al. (2024). We first give the path integral over $h({\bf r},t)$ and its dynamic conjugate field $\hat{h}({\bf r},t)$ Bausch et al. (1976); Täuber (2014) that is equivalent to and constructed from Eq. (2) of the main text together with the noise variance. The generating functional corresponding to Eq. (2) of the main text is given by Bausch et al. (1976); Täuber (2014)
@@
${\mathcal{S}}=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)\bigg{\}}.$ (23)
Using the one loop diagrams from the supplemental material of Ref. Jana et al. (2024) we obtain renormalized $D$ ($D^{<}$) and renormalized $\nu$ ($\nu^{<}$), we get
@@
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, where $S_{d}$ is the surface area of a $d$-dimensional unit sphere and $\gamma=\frac{\lambda_{1}}{\lambda}$ as defined in the main text. Using above equations and Setting $d=2$ in the above integrals, with $b=e^{\delta l}\approx 1+\delta l$ ($l$ is the RG time) and defining dimensionless coupling constant by $g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d}$ as in the main text, we obtain the RG flow equations for $D$ and $\nu$
@@
Furthermore, defining another dimensionless effective coupling constant, $g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}$ and using above equations we can write the flow equations of $g$ and $g_{2}$:
@@
$\frac{dg_{2}}{dl}$ $=$ $\frac{g_{2}}{2}[2-d-g\tilde{A}(\gamma)],$ (29)
@@
III MCT Calculations
We start with the general definitions of the correlation function $C({{\bf k}},\omega)$ and response function $G({{\bf k}},\omega)$:
@@ III MCT Calculations
The response and correlation functions are assumed to have the following scaling form in the long time large length scale limit:
@@
where $z$ is the dynamic exponent and $\chi$ is the roughness exponent; see also main text for their definitions. The Dyson’s equation for the self energy $\Sigma$ in the scaling limit is
@@
Here $\Gamma>0$ is the effective damping coefficient. The correlation function in Lorentzian approximation can be written as
@@
Figure 4(a) shows the one-loop diagrammatic correction to the self energy at ${\cal O}(\lambda^{2})$ that comes with a symmetry factor of 888 and contributes
[…]
After performing the $\Omega$-integral and making ${\bf q}\rightarrow{({\bf q}+\frac{\bf k}{2})}$, above integral becomes
$I$ $=\frac{\lambda^{2}D}{\Gamma}\int_{\bf q}\Bigl{(}{\bf q}+\frac{{\bf k}}{2}\Bigl{)}^{-2\chi-d+z}{\biggl{[}\Bigl{(}\frac{\bf k}{2}-{\bf q}\Bigl{)}\cdot\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}\biggl{]}}\frac{{\biggl{[}{\bf k}\cdot\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}\biggl{]}}}{{\Bigl{|}{\bf q}+\frac{\bf k}{2}\Bigl{|}}^{z}+{\Bigl{|}{\bf q}-\frac{\bf k}{2}\Bigl{|}}^{z}}$
$=\frac{\lambda^{2}D}{2\Gamma}\frac{2\chi+d}{2}\int_{\bf q}\frac{q^{2}({\bf k\cdot q})q^{-2\chi-d}}{q^{z}}\frac{({\bf k\cdot q})}{q^{2}}-\frac{\lambda^{2}D}{2\Gamma}k^{2}\int_{\bf q}\frac{q^{2}}{2q^{z}}q^{-2\chi-d}.$ (39)
@@
$\frac{\lambda^{2}D}{2\Gamma}\,\frac{2\chi+d}{2}\,\frac{k_{d}}{d}\times k^{2}\int_{\bf q}q^{-\chi-1}.$
@@
$\frac{\lambda^{2}D}{2\Gamma}\,\frac{k_{d}}{2}\,k^{2}\int_{\bf q}q^{-\chi-1}.$
@@
Figure 4(b) shows the one-loop contribution to the self energy at ${\cal O}(\lambda_{1}^{2})$ with a symmetry factor of 888 and contributes
[…]
$=\frac{4\lambda_{1}^{2}k_{i}k_{j}k_{n}}{k^{2}}\int_{{\bf q},\Omega}q_{i}q_{m}(k-q)_{j}(q-k)_{m}(q-k)_{n}\frac{1}{{\bf(q-k)^{2}}}\times\frac{2D\Gamma q^{-2\chi-d+z}}{(\Omega^{2}+\Gamma^{2}q^{2z})}\times\frac{1}{i\Omega+\Gamma{\bf|k-q|}^{z}}.$ (41)
After performing the $\Omega$-integral, we get
@@
Then shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, above integral becomes
$\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{n}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q-\frac{k}{2})_{m}(q-\frac{k}{2})_{n}(q+\frac{k}{2})_{m}}{{({\bf q}-\frac{{\bf k}}{2})}^{2}({|{\bf q}+\frac{\bf k}{2}|}^{z}+{|{\bf q}-\frac{\bf k}{2}|}^{z})}{|{\bf q}+\frac{\bf k}{2}|}^{-2\chi-d}.$ (43)
@@
$I_{2}$ $=\frac{4\lambda_{1}^{2}D}{\Gamma}\frac{k_{i}k_{j}k_{m}k_{n}}{k^{2}}\int_{\bf q}\frac{q_{i}q_{j}q_{m}q_{n}q^{-2\chi-d}}{2q^{2}q^{z}}\times\frac{2\chi+d}{2}=\frac{3\lambda_{1}^{2}D}{\Gamma d(d+2)}\frac{2\chi+d}{\chi}k_{d}k^{-\chi+2},$ (45)
where we have used the following identity Yakhot and Orszag (1986),
@@
Then, $I_{3}$ is evaluated by using the well-known relation Yakhot and Orszag (1986)
@@
Adding $I_{1}$, $I_{2}$ and $I_{3}$ we get contribution to zero frequency self energy at ${\cal O}(\lambda_{1}^{2})$:
@@
Figure 4(c) gives one of the two one-loop corrections to the self energy at ${\cal O}(\lambda\lambda_{1})$, with a contribution
@@
After performing the $\Omega$-integral and shifting ${\bf q}\rightarrow{\bf q}+\frac{\bf k}{2}$, we get
$\frac{2\lambda\lambda_{1}D}{\Gamma}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{{\bf q}}(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}\frac{|{\bf q}+\frac{\bf k}{2}|^{-2\chi-d}}{{|{\bf q}+\frac{\bf k}{2}|}^{z}+|\frac{\bf k}{2}-{\bf q}|^{z}}.$ (51)
@@
In evaluation of these integrals we have used identities (46) and (48). After some simplifications, the above integral gives
@@
After performing the $\Omega$-integral, above integral becomes